# MOON: Multi-Objective Optimization-Driven Object-Goal Navigation Using a Variable-Horizon Set-Orienteering Planner

### v2

@@ Abstract
Object-goal navigation (ON) enables autonomous robots to locate and reach user-specified objects in previously unknown environments, offering promising applications in domains such as assistive care and disaster response. Existing ON methods—including training-free approaches, reinforcement learning, and zero-shot planners—generally depend on active exploration to identify landmark objects (e.g., kitchens or desks), followed by navigation toward semantically related targets (e.g., a specific mug). However, these methods often lack strategic planning and do not adequately address trade-offs among multiple objectives. To overcome these challenges, we propose a novel framework that formulates ON as a multi-objective optimization problem (MOO), balancing frontier-based knowledge exploration with knowledge exploitation over previously observed landmarks; we call this framework MOON (MOO-driven ON). We implement a prototype MOON system that integrates three key components: (1) building on QOM [IROS05], a classical ON system that compactly and discriminatively encodes landmarks based on their semantic relevance to the target; (2) integrating StructNav [RSS23], a recently proposed training-free planner, to enhance the navigation pipeline; and (3) introducing a variable-horizon set orienteering problem formulation to enable global optimization over both exploration and exploitation strategies. This work represents an important first step toward developing globally optimized, next-generation object-goal navigation systems.
@@ I Introduction
In this work, we reformulate the standard single-target ON task as a multi-objective optimization-driven object goal navigation (MOON), and as a case study, model two conflicting objectives: frontier exploration for discovering unknown landmarks and efficient touring of previously observed landmarks (Fig. 2)111 The ”moon and star” representation was introduced as a schematic or cartoon-like abstraction to emphasize visual clarity while maintaining mathematical tractability. Nevertheless, our framework is not limited to point landmarks or circular sensing ranges and can be generalized to a wide range of application domains. . Our formulation emphasizes the intrinsic multi-objective nature of ON tasks and is both orthogonal and complementary to existing problem settings, such as multi-ON tasks [10], which address multiple target objects in an optionally specified order, and general multi-objective navigation [11]. Specifically, we model ON planning as a variable-horizon optimization problem aiming to maximize average SPL (Success weighted by Path Length [12]) based on incrementally constructed landmark maps. Although our method is compatible with the TFP, RLP, and ZSP frameworks, we adopt the TFP framework from [1] as the basis for our experiments. In this framework, the robot detects objects from RGBD and semantic images and semantically estimates landmark objects related to the target. We then extend TFP by introducing a query-based occupancy map (QOM) [13], which compactly describes essential landmark information by discarding data irrelevant to object-target navigation and focusing solely on relevant landmarks. Finally, we formulate the landmark visiting order optimization problem (LOOP) as a set orienteering problem (SOP) and further extend it to a variable-horizon SOP that incorporates map updates. Notably, the proposed method facilitates map sharing and multi-objective navigation, enhancing the efficiency and scalability of ON tasks by integrating semantic reasoning with optimal planning.
@@ III A MOON Planner
subject $\text{subject to}\quad\mathbf{x}\in\mathcal{X}$

### v3

@@ Abstract
This paper proposes MOON (Multi-Objective Optimization-driven Object-goal Navigation), a novel framework designed for efficient navigation in large-scale, complex indoor environments. While existing methods often rely on local heuristics, they frequently fail to address the strategic trade-offs between competing objectives in vast areas. To overcome this, we formulate the task as a multi-objective optimization problem (MOO) that balances frontier-based exploration with the exploitation of observed landmarks. Our prototype integrates three key pillars: (1) QOM [IROS05] for discriminative landmark encoding; (2) StructNav [RSS23] to enhance the navigation pipeline; and (3) a variable-horizon Set Orienteering Problem (SOP) formulation for globally coherent planning. To further support the framework’s scalability, we provide a detailed theoretical foundation for the budget-constrained SOP formulation and the data-driven mode-switching strategy that enables long-horizon resource allocation. Additionally, we introduce a high-speed neural planner that distills the expert solver into a transformer-based model, reducing decision latency by a factor of nearly 10 while maintaining high planning quality.
@@ I Introduction
In this work, we reformulate the standard single-target ON task as a multi-objective optimization-driven object goal navigation (MOON), and as a case study, model two conflicting objectives: frontier exploration for discovering unknown landmarks and efficient touring of previously observed landmarks (Fig. 2) [^1] The ”moon and star” representation was introduced as a schematic or cartoon-like abstraction to emphasize visual clarity while maintaining mathematical tractability. Nevertheless, our framework is not limited to point landmarks or circular sensing ranges and can be generalized to a wide range of application domains. . Our formulation emphasizes the intrinsic multi-objective nature of ON tasks and is both orthogonal and complementary to existing problem settings, such as multi-ON tasks [10], which address multiple target objects in an optionally specified order, and general multi-objective navigation [11]. Specifically, we model ON planning as a variable-horizon optimization problem aiming to maximize average SPL (Success weighted by Path Length [12]) based on incrementally constructed landmark maps. Although our method is compatible with the TFP, RLP, and ZSP frameworks, we adopt the TFP framework from [1] as the basis for our experiments. In this framework, the robot detects objects from RGBD and semantic images and semantically estimates landmark objects related to the target. We then extend TFP by introducing a query-based occupancy map (QOM) [13], which compactly describes essential landmark information by discarding data irrelevant to object-target navigation and focusing solely on relevant landmarks. Finally, we formulate the landmark visiting order optimization problem (LOOP) as a set orienteering problem (SOP) and further extend it to a variable-horizon SOP that incorporates map updates. Notably, the proposed method facilitates map sharing and multi-objective navigation, enhancing the efficiency and scalability of ON tasks by integrating semantic reasoning with optimal planning.
@@ III A MOON Planner
$\text{subject to}\quad\mathbf{x}\in\mathcal{X}$